OpenStudy (anonymous):

If ratios m/5, 14/10, and 21/n are proportional, then which is proportional to m+n/n-m?

OpenStudy (anonymous):

A) 13/4 B) 13/5 C) 11/4 D) 11/5

OpenStudy (anonymous):

m=7 n=15 7+15/15-7 = 11/4 = option C

OpenStudy (anonymous):

Thank you!